8 A ball is kicked at an angle ? = 45°. It is intended that the ball lands in the back of a moving truck which has a trunk of length L = 2.5 m. If the initial horizontal distance from the back of the truck to the ball, at the instant of the kick, is do = 5 m, and the truck moves directly away from the ball at velocity V = 9 m/s (as shown), what is the maximum and minimum velocity vo so that the ball lands in the trunk. Assume that the initial height of the ball is equal to the height of the ball at the instant it begins to enter the trunk.

8



A ball is kicked at an angle ? = 45°. It is intended that the ball lands in the back of a moving truck which has a trunk of length L = 2.5 m. If the initial horizontal distance from the back of the truck to the ball, at the instant of the kick, is do = 5 m, and the truck moves directly away from the ball at velocity V = 9 m/s (as shown), what is the maximum and minimum velocity vo so that the ball lands in the trunk. Assume that the initial height of the ball is equal to the height of the ball at the instant it begins to enter the trunk.

Grade:12

1 Answers

Saurabh Singh
10 years ago
speed of ball relative to truck Vr= VoCos 45 -9m/s
time taken by ball T= (VoSin45/9.8)*2

horizontal distance travelled by ball relative to truck H = Vr/T

MAximum distance that the ball can travel = 5+2.5 = 7.5m
Minimum distnace the ball can travel = 5m

hence 5<=H<=7.5

we can substitute the values and get the answer

Thanks & Regards

Saurabh Singh,

askIITians Faculty

B.Tech.

IIT Kanpur